OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (0, −4) and a directrix of y = 4... I don't understand this at all please help :) ( not just the answer the concept)

OpenStudy (jdoe0001):

in a parabola, it's focus and directrix are equidistant from the vertex of the parabola meaning, from the vertex to the focus, is the same distance as from the vertex to the directrix meaning that the vertex is half-way in between those two fellows so what are the coordinates for the vertex you'd think?

OpenStudy (anonymous):

Is the vertex 0 ?

OpenStudy (jdoe0001):

well, the vertex has an x-coordinate and y-coordinate.... so.... what do you think?

OpenStudy (anonymous):

(0,0)

OpenStudy (anonymous):

also i know something has to do with the equation 1/4p= a ....

OpenStudy (jdoe0001):

well yes, the vertex is at (0, 0) also keep in mind that the parabola opens up towards the focus and away from the directrix so  meaning is vertical and is going downwards so \(\bf (x-h)^2=4p(y-k) \quad p=\textit{distance from the vertex to the focus}\) since the parabola is going downwards, that means that "p" will be negative